Calculate the partial pressure (in atm) of Ar in a flask that contains a mixture of mols 1. _____ of Ar and 5.0 mols of Ne in the flask with a total pressure of 2.6 atm.


1
Expert's answer
2021-12-06T16:11:02-0500

"P_{Ar}=\Chi_{Ar}P_{Total}\\\Chi_{Ar}=\frac{Moles\>of\>Ar}{Total Moles}" "=\frac{1}{6}"


P_{Ar}="\frac{1}{6}×2.6=0.4333\>Atm"